How to Synthesis 2-Butene to get tenormin?
Answer #1

The Wittig reaction or Wittig olefination is a chemical reaction of an aldehyde or ketone with a triphenyl phosphonium ylide (often called a Wittig reagent) to give an alkene.

R1 Ph3-P+ C=O + C-R3 — R1 R3 C=c R2 R4 + Ph3P=0 R2 R4